OpenStudy (anonymous):

If Elaine scored an 80 on her math test and predicted her score with 15% error, what are the two possible predictions that Elaine could have had for her math test?

OpenStudy (anonymous):

PLEASE HELP!!!! I WILL MEDAL!!!!!!!

OpenStudy (anonymous):

@Luigi0210 , @Directrix , @paki

OpenStudy (anonymous):

I have limited time so please help me!!!!

OpenStudy (inowalst):

Is it a test?

OpenStudy (anonymous):

no but I still have only a few minutes to complete it

OpenStudy (inowalst):

<.< Then what is it?

OpenStudy (anonymous):

It is for my DBA with my teacher which is a Discussion based assignment that I have to do in order to finish my module

OpenStudy (anonymous):

in math

OpenStudy (inowalst):

I know what a DBA is and if its for a DBA its for an oral exam aka a DBA.

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

I need help so please help I will fan and medal!! for whoever answers correctly

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@mathstudent55

OpenStudy (ribhu):

69.56 & 94.12

OpenStudy (ribhu):

your answers.

OpenStudy (anonymous):

both of them are answers?

OpenStudy (anonymous):

I have got a couple of other questions.

OpenStudy (ribhu):

k

OpenStudy (anonymous):

For each equation below, determine whether the highlighted number is the part, the whole, or the percent. A. 80=0.6*b B. a=0.35*75 C. 90=32%*b D. 15=p%*70 (highlighted numbers are: 0.6, 75, 90, 70)

OpenStudy (ribhu):

question is not very clear

OpenStudy (ribhu):

i really didn't get it if u cud make me understand then i can do it

OpenStudy (anonymous):

so for each equation find the highlighted numbers and determine if that number is the part, the whole, or the percent.

OpenStudy (ribhu):

still @mjperez18 get to another ques

OpenStudy (anonymous):

so are the numbers 0.6,75,70,and 90 the part, whole, or percent of the equation they are in

OpenStudy (ribhu):

i got it probably

OpenStudy (anonymous):

k

OpenStudy (ribhu):

A is part B is the whole.

OpenStudy (ribhu):

C is the whole & D is the whole.

OpenStudy (anonymous):

. Find the solution to each proportion problem below: a. x/15=12/5

OpenStudy (anonymous):

b. 20/x = 35/7

OpenStudy (anonymous):

c. 3/10 = x/36

OpenStudy (anonymous):

d. 81/27 = 18/x

OpenStudy (ribhu):

a = 36

OpenStudy (anonymous):

Group the ratios below together by placing equivalent ratios in the same circle. 45% 15:40 12 to 32 4:10 6 to 8 27 to 60 20 to 50 4

OpenStudy (ribhu):

next

OpenStudy (anonymous):

okay but csn you help me with this one later?

OpenStudy (ribhu):

which one?

OpenStudy (anonymous):

The one I just asked

OpenStudy (anonymous):

I think the question means pair two ratios together that are equivalent

OpenStudy (anonymous):

I just looked at my paper and that is the last question

OpenStudy (anonymous):

@ribhu

OpenStudy (anonymous):

Group the ratios below together by placing equivalent ratios in the same circle. 45%, 15:40, 12 to 32, 4:10, 6 to 8, 27 to 60, 20 to 50, 4

OpenStudy (ribhu):

i told u this, 27:60

OpenStudy (anonymous):

but you have to pair two ratios together

OpenStudy (anonymous):

so say 24:43 is equal to 8:43 that would be a pair

OpenStudy (anonymous):

does that make sense @ribhu

OpenStudy (ribhu):

45% with 27:60

OpenStudy (anonymous):

k

OpenStudy (ribhu):

15:40 with 12to 32

OpenStudy (ribhu):

4:10 with 20 to 50

OpenStudy (ribhu):

6/8 doesn't match with 4 pls see

OpenStudy (anonymous):

do either of those match any of the others.

OpenStudy (ribhu):

doubt with the last all others are properly matched.

OpenStudy (anonymous):

could you check please?

OpenStudy (ribhu):

i think i have checked for this one.

OpenStudy (ribhu):

did u copy the question correctly

OpenStudy (anonymous):

would it make a difference if i changed one to 20:54? other than that there is nothing else there.

OpenStudy (ribhu):

no something else

OpenStudy (anonymous):

not all of them copied

OpenStudy (ribhu):

then leave this and check rest

OpenStudy (anonymous):

9/20, 45%, 15:40, 10/25, 12 to 32, 4:10, 15/20, 6 to 8, 27 to 60, 3/8, 20:50, 45:60

OpenStudy (ribhu):

9/20 =45%

OpenStudy (ribhu):

15/40 = 3/8

OpenStudy (ribhu):

10/25= 4/10

OpenStudy (ribhu):

12/32 = 3/8

OpenStudy (anonymous):

there should be 4 answers using all 12 of the ratios because I am only given 4 circles to put them in. So three in each pair or triplet

OpenStudy (ribhu):

15/20 = 6/8

OpenStudy (ribhu):

now u can match up??

OpenStudy (anonymous):

thats five answers I only need 4 but with three in each group

OpenStudy (ribhu):

9/20 45 27 to 60

OpenStudy (ribhu):

now u group

OpenStudy (ribhu):

15/40 12/32 3/8

OpenStudy (ribhu):

10/25 4/10 20/50

OpenStudy (ribhu):

15/20 6/8 45/60

OpenStudy (ribhu):

@mjperez18 i think this solves it

OpenStudy (anonymous):

thank you